Gilles Perrot is a Ph.D student at the university of Franche-comté within the FEMTO-ST Laboratory. His research interests currently focus on GPU-oriented, high performance image processing algorithms. His university studies have been conducted at University of Orsay (Paris XI) and at École Normale Supérieure (ENS). He received the M.Sc. degree in electrical engineering in 1993. He received the Agrégation of electrical engineering (EE) in 1994 (the highest qualification of teachers in France). In 1995 he presented a MA thesis in signal processing. From 1995 to 2009, he has been a EE teacher, technical advisor or manager for several projetcs on Communication and Information technologies.

Stéphane Domas is Assistant Professor at the University of Franche-Comté. He is a member of the AND team, within the DISC department of the FEMTO-ST Laboratory. His domain of research is the high performance computing applied to a large range of architectures, from FPGAs and GPUs to widely distributed grid of machines/clusters. More precisely, he is concerned about the development of programming/execution environments for such architectures, and the creation or the optimization of algorithms that take into account their hardware specificities.

Raphaël Couturier received the Ph.D. degree in 2000 in Computer Science from the Henri Poincar University in Nancy, France. From 2000 to 2006 he was an Assistant Professor in the University of Franche Comte. Then he has been a Professor in the same University. His research interests include parallel and distributed algorithms with a strong knowledge on asynchronous iterative methods, and GPU and FPGA computing. Raphal authored or co-authored more than 70 papers in conferences and journals and a book. He has also served in many program committees for conferences.

Nicolas Bertaux graduated from École Normale Supérieure (ENS). He received the M.Sc. degree in electrical engineering from Paris University, Orsay, France in 1993 and the Agrégation of electrical engineering (EE) in 1994 (the highest qualification of teachers in France). He received the Ph.D. from ENS in january 2000 in electrical engineering. He is an assistant professor in leectrical engineering at École Centrale de Marseille. His main research interests at Fresnel Institute (UMR CNRS 7249) include digital image processing and signal processing.